OpenStudy (anonymous):

Find the quotient of quantity of 8 times x to the 2nd power times y plus 12 times x times y to the 3rd power minus 16 times x times y all over 4 times x times y. 4x + 8y2 − 12 12x2 − 16x + 20 2x3y2 + 3x2y4 − 4x2y2 2x + 3y2 − 4

OpenStudy (gw2011):

The answer is: 2x+3y^2-4 which is the last choice given The simplest way of solving this problem is by writing it as: 8x^2y/4xy + 12xy^3/4xy - 16xy/4xy Divide each part and you get: 8x^2y/4xy = 2x 12xy^3/4xy = 3y^2 -16xy/4xy = -4